Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-139_2001-145

Mission Operations Days: 2001/139 to 2001/145
May 19, 2001 00:00:00 GMT to May 25, 2001 00:00:00 GMT


NOTE: On day 140 at 20:48:26z, during a MODIS science buffer playback, there were 364,464 uncorrectable CADU's received by EDOS. This has happened many times since then, each having an identical amount of CADU's as uncorrectable. It has been determined that the uncorrectables have originated inside supersets 30 and 31 in PWA 16 in the SSR. These two supersets have been moved to the FREE list in the SSR and MODIS is currently (as of day 145) collecting data with only 33 supersets. The course of action by both the FOT and the MODIS science team is still undecided.

EDOS is also receiving repeated packets at the same time as the uncorrectables and this is causing them not to be able to process any of the good MODIS science data as of yet. Furthermore, with the reduced buffer capacity, and with no change in the data collection rate, there is a strong probability that there will be continued data losses to MODIS due to lock up losses.

MODIS continues to operate nominally.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode with SSR supersets 30 and 31 unallocated
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)Nominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off; #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 2 set of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 5 Gain Change (20 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68Nominal

This Week's Completed MODIS Activities:

Saturday, May 19, 2001

None

Sunday, May 20, 2001

None

Monday, May 21, 2001

None

Tuesday, May 22, 2001

None

Wednesday, May 23, 2001

2001/143 18:03:00 Inhibit ID's 53, 33, and 58 are set. (In response to SSR Anomaly)

Thursday, May 24, 2001

2001/144 14:30:00 Inhibit ID's 53, 33, and 58 are cleared. (Set from earlier SSR Anomaly)

Friday, May 25, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

Since the Solar Diffuser door movement is a limited life item, and the data from the MODIS playbacks may be lost, there were no SD/SDSM activities performed this week.

143/18:03:00 As a result of the SSR anomaly, Inhibit ID's 53 (SD/SDSM Commands), 33 (CP Macro Commands), and 58 (Door Motor Power Commands) are set to inhibit the following SD/SDSM activities from happening:
143/00:49:57 ATC OA16: SDSM SCREENED
143/02:28:50 ATC OA15: SDSM OPEN

Upcoming MODIS Events:

Saturday, May 26, 2001

None

Sunday, May 27, 2001

None

Monday, May 28, 2001

None

Tuesday, May 29, 2001

None

Wednesday, May 30, 2001

None

Thursday, May 31, 2001

None

Friday, June 01, 2001

None

Maneuvers:

There will be no MODIS Lunar Roll Maneuver in May because the roll angle would have been too great (more than 20-degree roll required).

The next MODIS Lunar Roll Maneuver is scheduled for June 10, 2001.

MODIS Anomalies:

None

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Telemetry is trending nominally.

Non-MODIS Significant Events:

MOPITT Anomaly:

MOPITT's Displacement Side B position went Red (at 2001/127-04:07) during a Real-Time contact on DOY 127. The Flight Controllers sent the Red Limit Response procedure to Safe the Instrument (at 2001/127-04:18).

After analyzing telemetry history, the MOPITT team has determined the event was not critical. MOPITT has been commanded into its Standby Mode. The MOPITT IOT plans to transition out of Standby to Science Mode sometime next week pending final analysis verification.

Limited Life Item Status:

SRCA 10W Lamp #1: 186.0 of 500 hours
SRCA 10W Lamp #2: 142.9 of 500 hours
SRCA 10W Lamp #3: 152.1 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.5 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1615 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements